question: Five more than six times a number is eleven. Find the number.
I need helping writing and solving the equation

Found 2 solutions by ikleyn, ankor@dixie-net.com:
Answer by ikleyn(52777)

.

Let x be the number (now unknown).


Then "six times the number" is 6x.


"Five more than six times a number" is 6x + 5.


"Five more than six times a number is eleven" means that


    6x + 5 = 11.


This equation is exact translation of English to Math.


To solve the equation, subtract 5 from both sides.  You will get


    6x = 11 - 5

    6x = 6.


Now divide both sides by 6.  You will get


    x   = 1.


ANSWER.  The number is 1.

Solved, explained and completed.



Answer by ankor@dixie-net.com(22740)

Write an equation for exactly what it says
:
Five more than six times a number is eleven.
6x + 5 = 11
subtract 5 from both sides
6x = 11 - 5
6x = 6
divide both sides by 6
x = 1 is the number.
:
Check solution x=1 in original equation
6(1) + 5 = 11
6 + 5 = 11